# Varying Newton constant,entropy and the black hole evaporation law

### v1

@@ 4 Compact systems with an event horizon
using $\sigma^{0}=u^{0}\sigma=|g_{00}|^{-\frac{1}{2}}\sigma$ and assuming that $\beta$ depends only on time we can rewrite eq.(29) as
$\begin{array}[]{l}\beta\partial_{t}G\int d{\bf x}\sqrt{\gamma}\rho+G\partial_{t}S+\frac{\beta}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda=0.\end{array}$ (31)
We define the mass
$M=\int d{\bf x}\sqrt{\gamma}\rho.$ (32)
We skip the time derivative of $\Lambda$ in eq.(31)for an application to the black hole. Then, eq.(31) can be rewritten as
@@
We intend to apply the formulae (2) and (24) to gravitational systems with an event horizon whose entropy and temperature has been determined by other methods [40][41][42]. In classical closed systems the total entropy does not change ( neither does it in quantum systems). In systems with an event horizon we may restrict ourselves to subsystems located inside or outside the horizon. Then, their entropy may depend on time. The corresponding calculations have been performed in quantum systems ( calculation of the entanglement entropy) [43][44][45]( see the reviews in [46][47]). We think that this entropy can be incorporated in eq.(33). Eq.(33) is a consequence of eq.(24). In eq.(24) we have replaced the covariant derivative of the energy-momentum tensor (the lhs of eq.(18)) appearing in the Bianchi identity (23) by the covariant derivative of the entropy density according to the first law of thermodynamics (the rhs of eq. (18)). The quantum results for the entropy and for the temperature (obtained by calculations of the partition function [43] or from the quantum gravity in the Euclidean formulation [42]) show that $S\simeq\frac{1}{\hbar}$ and $T$ is proportional to $\hbar$. Hence, their product on the rhs of eqs.(27) and (33) has a classical meaning.
The formula for the temperature of the Schwarzschild black hole of mass $M$ is [41]
$T_{BH}=\frac{\hbar}{8\pi GM}.$ (34)
It can be derived from the correlation functions of the quantum scalar fields on the Schwarzschild background or from the Rindler space approximation on the basis of the KMS periodicity argument. Another derivation is based on Euclidean quantum gravity in one-loop approximation [42]. The entropy of the Schwarzschild black hole can be calculated from the second law of thermodynamics $dM=T_{BH}dS_{BH}$ on the basis of eq.(34)
@@
(it is also obtained from a path integral as an entanglement entropy [42][43][44][45], in such a case it would appear in the semi-classical expansion of the effective action). We insert the black hole entropy (35) in eq.(33). Then, we obtain
$\frac{3}{2}\frac{d\ln G}{dt}=-\frac{d\ln M}{dt}.$ (36)
Hence,
$G=G_{0}M_{0}^{\frac{2}{3}}M^{-\frac{2}{3}}.$ (37)
The change of mass $M$ resulting from the black hole evaporation through an emission of massless particles is described by the formula [48] [49][50][51] (see also the reviews in [46][47])
@@
where $\alpha$ is a number which can be obtained from Stefan-Boltzmann formula multiplied by the black hole surface and an estimate of emissivity [48] ( the formula does not apply if $T_{BH}$ is lower than the temperature of the environment and for a small time). Eqs.(38) and (33) express the thermodynamic conservation law. The black hole is releasing the energy and the entropy gained during a mass accretion in the process of black hole formation.
Eq.(38) has the solution
$M^{3}=M_{0}^{3}-3\hbar\alpha\int_{t_{0}}^{t}G^{-2}(s)ds,$ (39)
where $t_{0}$ is the time of the black hole formation. We can insert this $M$ in eq.(37) for $G$. Denote
$X=\int_{t_{0}}^{t}G^{-2}(s)ds.$ (40)
Then, eq.(37) can be rewritten as
$\frac{dX}{dt}=G_{0}^{-2}M_{0}^{-\frac{4}{3}}(M_{0}^{3}-3\hbar\alpha X)^{\frac{4}{9}}.$ (41)
The resulting equation for $G$ has the solution
$G=G_{0}\Big(1-\frac{5}{3}\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{-\frac{2}{5}}.$ (42)
Inserting $G$ of eq.(42) into the formula (39) for the mass we obtain
$M=M_{0}\Big(1-\frac{5}{3}\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{\frac{3}{5}}\equiv M_{0}Z^{\frac{3}{5}}.$ (43)
It follows that
$\begin{array}[]{l}T_{BH}=\frac{\hbar}{8\pi G_{0}M_{0}}Z^{-\frac{1}{5}}\end{array}$ (44)
The entropy tends to zero as $GM^{2}\simeq Z^{\frac{4}{5}}$, the luminosity $I$ of the black hole is proportional to $T_{BH}^{4}$ and to the area of the black hole ($A\simeq(GM)^{2}$ ). Hence, the luminosity behaves as
$I\simeq Z^{-\frac{2}{5}}$ (45)
tending to infinity at the end of the black hole’s life time with a speed lower than the one for a fixed Newton constant.
The black hole’s radius $2GM$ tends to zero as $Z^{\frac{1}{5}}$ but the average energy density defined as the mass divided by the volume remains constant as
$M\Big(\frac{4\pi}{3}(2GM)^{3}\Big)^{-1}=M_{0}\Big(\frac{4\pi}{3}(2G_{0}M_{0})^{3}\Big)^{-1}$ (46)
@@
This is different than in the Hawking-Page model [52][51] when the averaged energy density tends to infinity at the end of black hole’s life time. The results (42)-(46) could be compared with observations of the gamma-ray bursts [24]. In particular, with the recent observations of the giant neutrino emissions [23] which by some authors [21][22] are interpreted as resulting from primordial black hole explosions.
The life time of the black hole according to eq.(43) is
$t_{f}=\frac{3}{5}G_{0}^{2}M_{0}^{3}\hbar^{-1}\alpha^{-1}.$ (47)
It is longer by a factor $\frac{9}{5}$ than the one of the Hawking-Page model (which would result from the assumption that $G=const$ in eq.(39)).
5 Summary and outlook
We have shown that the Bianchi identity of classical Einstein equations admits a violation of the ideal fluid energy-momentum conservation if the Newton “constant” G and the cosmological term depend on time. As a generalization of the first and the second laws of thermodynamics the covariant energy-momentum divergence can be expressed by the covariant divergence of the entropy density. Then, the entropy density enters the Bianchi identity. In this paper we have applied the formula for the change of entropy during the black hole evaporation to determine the time-dependence of $G$ in the Bianchi identity. As a consequence of a growing $G$ the black hole evaporation is slower and the average black hole density is constant in contradistinction to the one for a fixed $G$ (this could be subject to astronomical observations) . The time-dependence of $G$ is unavoidable if we wish to describe the non-conserved energy-momentum of an evaporating black hole by Einstein equations (2). The increasing $G$ could have a physical stabilizing origin preventing the appearance of an infinite density when the black hole’s radius tends to zero. Another problem deserving a study concerns the entropy of an expanding universe with a cosmological horizon. For such a system the varying cosmological constant in the Bianchi identity should be relevant.
@@ 5 Summary and outlook
Acknowledgement A part of the results of this paper is contained in JH’s bachelor thesis. JH thanks Dr. Aleksander Kozak for supervising the thesis.

### v2

@@ 4 Compact systems
using $\sigma^{0}=u^{0}\sigma=|g_{00}|^{-\frac{1}{2}}\sigma$ and assuming that $\beta$ depends only on time we can rewrite eq.(37) in the form (25)
$\begin{array}[]{l}\beta\partial_{t}G\int d{\bf x}\sqrt{\gamma}T^{0}_{0}+G\partial_{t}S+\frac{\beta}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda=0.\end{array}$ (40)
When we skip the time derivative of $\Lambda$ in eq.(40)and denote $M=\int d{\bf x}\sqrt{\gamma}T^{0}_{0}$ (we admit that with $\Sigma^{0}_{0}\neq 0$ the spatial integrals over $\rho$ and $T^{0}_{0}$ may be different) then we obtain
@@
5 Mass loss in evaporation
In this section we discuss dynamics of a change of mass of a compact object. We assume that such dynamics can be described in the framework of classical Einstein gravity if $G$ and $\Lambda$ are allowed to change in time. We have two equations (24) and (25) for a relation between time derivatives of $G$ and $M$. We have in mind a dynamical system resulting from a star contraction which subsequently is decaying through an emission of radiation. An investigation of the dynamics of a star collapse to the black hole has been initiated by Oppenheimer and Snyder [49]. Then, Einstein and Straus [50] [51] studied a star (Schwarzschild metric) in an expanding universe by pasting the FLWR metric with the Schwarzschild solution. In this framework there is a lot of work recently on the formation of the black hole [52] [53], its evolution in the FLWR space-time [34][35][36][37] and final evaporation [54][55] (nonsingular black holes). In this section we apply eqs.(25) and (27) to a description of a decrease of mass as a result of an emission of radiation of a compact object (a black hole) with a hot event horizon.
A change of mass $M$ resulting from an emission of radiation from a surface can be obtained from the Stefan-Boltzmann law (we assume that the temperature $T$ is larger then the temperature of the environment and that the absorption from the environment can be neglected)
$\frac{dM}{dt}=-\lambda T^{4}A,$ (42)
where $\lambda$ is the Stefan-Boltzmann constant ($\lambda\simeq\hbar^{-3}$) and $A$ is the area of the surface. For a compact system with an event horizon we can obtain a system of equations (24)-(25) and (42) depending on $G$ and $M$ because the area and temperature are functions of $G$ and $M$. In eq.(24) we would need to know the pressure whereas in eq.(25) the entropy as functions of $G$ and $M$ in order to determine a relation between these variables. We consider a power-law Ansatz as a possible solution of eqs.(24)-(25)
$G=G_{0}M_{0}^{\gamma}M^{-\gamma}.$ (43)
It can be seen that if $\partial_{t}\Lambda=0$ and $p=0$ then from eq.(24) $\gamma=1$. In order to determine a relation between $G$ and $M$ from eq.(25) we need to know the temperature $T$ and entropy $S$. We consider a simple power-law Ansatz
$T=\theta G^{\alpha_{1}}M^{\alpha_{2}},$ (44)
$S=\kappa G^{\nu_{1}}M^{\nu_{2}}.$ (45)
We substitute it in eq.(25). Then, we obtain
$\gamma=\theta\kappa\nu_{2}(1+\theta\kappa\nu_{1})^{-1}$ (46)
with $\alpha_{1}=-\nu_{1}$$\alpha_{2}=1-\nu_{2}$. If in addition we demand the definition of temperature as
$\frac{1}{T}=\frac{\partial S}{\partial M}$ (47)
then we get the relation $\kappa\theta\nu_{2}=1$ and $\gamma=\nu_{2}(\nu_{1}+\nu_{2})^{-1}$.
As conjectured by Bekenstein [56] the entropy of the black hole is proportional to its area. The calculation of the temperature of the black hole in a system of quantum scalar particles on a Schwarzschild background [57] gave the result (the formula has been established for various quantum systems defined on the Schwarzschild background [58][59])
$T_{BH}=\frac{\hbar}{8\pi GM}$ (48)
Then, from the relations (44)-(47) we obtain the entropy
@@ 5 Mass loss in evaporation
If instead of (47) we use (48) and
$dS=T_{BH}^{-1}dM$ (50)
appearing in the black hole mechanics in [44] then the formula for the entropy is
$S=\frac{8\pi}{\hbar}\int GMdM.$ (51)
For a time-dependent $G$ eqs.(49) and (51) are different. We discuss all options (49), (51)and (44)-(45) leading to the relation (43) with different $\gamma$. The reason is that a time-dependent $G$ may appear in an effective field theory of various origin [60]: Higgs gauge theories with symmetry breaking [61][62],Branse-Dicke [10][11][63], $f(R)$ and higher derivatives gravities [64][63]. For a constant $G$ the entropy (49) has been derived as an entanglement entropy in various models of quantum field theory [58][60][65][66]. A time dependent temperature (possibly in eq.(51)) can appear if the black hole surface is time dependent ( a trapping horizon) in an expanding universe [67]. A renormalization of the Newton constant as in refs. [13][14][15][16] can lead to a modified formula for the entropy. A microscopic counting of states in loop quantum gravity [68] does not fix the coefficient for entropy in front of the area as well. We consider also the option
$S_{\kappa}=\frac{\kappa}{4G\hbar}A,$ (52)
where $\kappa=G_{eff}^{-1}G$ with an effective coupling $G_{eff}$ which can come from a calculation of the entropy from the partition function. With the entropy (52) we get the relation (43) with $\gamma=\frac{\kappa}{1+\frac{1}{2}\kappa}$ .
Summarizing, $\gamma=1$ for zero pressure in eq.(25) or the entropy (51), $\gamma=\frac{2}{3}$ for the Bekenstein-Hawking entropy (49). The remaining values of $\gamma\geq 0$ are rather hypothetical. However, they lead to interesting consequences. A constant $G$ (corresponding to $\gamma=0$ ) cannot describe evaporation by means of Einstein equations because owing to Bianchi identities decreasing $M$ (with constant $\Lambda$) requires a change of $G$.
We assume in all models $T\simeq(GM)^{-1}$ and $A\simeq(GM)^{2}$. When we insert the temperature and the area in eq.(42) then we obtain [69] [70][71][72] (see also the reviews in [73][74][75])
@@
where $\alpha$ is a ($\hbar$ independent) number which can be obtained from the Stefan-Boltzmann constant and an estimate of emissivity [69].
If the pressure is zero in eq.(24) (as in the Oppenheimer-Snyder approximation) leading to eq.(27) and $\partial_{t}\Lambda$ is negligible in eq.(27) then (as $GM\simeq const$) we obtain from eq.(53)
$M_{t}=M_{0}-\hbar\alpha M_{0}^{-2}G_{0}^{-2}(t-t_{0}),$ (54)
where $t_{0}$ is the time of the black hole formation. We obtain the same result ($\gamma=1$) from eq.(25) with an application of the formulas (50)-(51) for the entropy.
With a varying $\Lambda$ and $\gamma=1$ we substitute from eq.(27) in eq.(53)
$\begin{array}[]{l}MG=G_{0}M_{0}-\int_{t_{0}}^{t}\frac{ds}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{s}\Lambda.\end{array}$ (55)
Then,
$M=M_{0}-\hbar\alpha\int_{t_{0}}^{t}d\tau\Big(G_{0}M_{0}-\int_{t_{0}}^{\tau}\frac{ds}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{s}\Lambda\Big)^{-2}$ (56)
Note that from eq.(55)it follows that if $\partial_{t}\Lambda=0$ then the temperature of the black hole as well as its luminosity are constant during the whole black hole’s life time (there is no black hole explosion, contrary to [72]). With the varying cosmological term as a consequence of eq.(55) the variation of the temperature and luminosity of the black hole depends on the sign of $\partial_{t}\Lambda$ .
If eqs.(48)-(49) are applied in eq.(25) then
$\begin{array}[]{l}\frac{3}{2}M\partial_{t}G+G\partial_{t}M+\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda=0.\end{array}$ (57)
If $\partial_{t}\Lambda=0$ then from eq.(57) we obtain $\gamma=\frac{2}{3}$. Surprisingly, this is the same relation between the Newton constant and the mass which follows from the formula for the mass of the neutron star and its Chandrasekhar limit [45] (sec.11) According to ref.[18] it could serve as an estimate of the variation of the Newton constant.
Let us consider the general option (43) with $\gamma\neq 0$ . When we substitute the relation (43) between $G$ and $M$ in eq.(53) then we obtain the solution for $M$ ( when $\gamma\neq\frac{3}{2}$)
$M=M_{0}\Big(1-(3-2\gamma)\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{\frac{1}{3-2\gamma}}\equiv M_{0}Z^{\frac{1}{3-2\gamma}},$ (58)
$G=G_{0}\Big(1-(3-2\gamma)\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{-\frac{\gamma}{3-2\gamma}}\equiv G_{0}Z^{-\frac{\gamma}{3-2\gamma}}.$ (59)
When $\gamma=\frac{3}{2}$ then
$M=M_{0}\exp\Big(-\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)$ (60)
It follows that ( for $\gamma\neq\frac{3}{2}$)
$\begin{array}[]{l}GM=G_{0}M_{0}Z^{\frac{1-\gamma}{3-2\gamma}}\end{array}$ (61)
$\begin{array}[]{l}GM^{2}=G_{0}M_{0}^{2}Z^{\frac{2-\gamma}{3-2\gamma}}\end{array}$ (62)
and
$\begin{array}[]{l}G^{-3}M^{-2}=G_{0}^{-3}M_{0}^{-2}Z^{\frac{3\gamma-2}{3-2\gamma}}\end{array}$ (63)
These equations have some consequences for the behaviour of the temperature, luminosity (L), entropy and the mean density (defined as the mass divided by the volume). Clearly, $M$ tends to zero for all $\gamma$ and $G$ tends to infinity. However, the dependence of the temperature, luminosity, entropy and the mean desity may substantially depend on $\gamma$. For the temperature and luminosity we have
$\gamma$ T,L time
$0<\gamma<1$ increasing to infinity finite
$\gamma=1$ constant all time
$1<\gamma<\frac{3}{2}$ decreasing to zero finite
$\gamma\geq\frac{3}{2}$ decreasing to zero infinite
For $\gamma\geq\frac{3}{2}$ the mass, the temperature and the luminosity tend to zero asymptotically in time.
@@
Let us note that the entropy (62) is decreasing to zero if $\gamma<2$, is constant if $\gamma=2$ and is growing in time if $\gamma>2$. The mean density (63)is tending to infinity if $\gamma<\frac{2}{3}$, is constant for $\gamma=\frac{2}{3}$ and is tending to zero if $\gamma>\frac{2}{3}$ (this is different than in the Hawking-Page model [72]). The results could be compared with observations of the gamma-ray bursts [26]. In particular, with the recent observations of the giant neutrino emissions [25] which by some authors [23][24] are interpreted as resulting from primordial black hole explosions.
The life time of the black hole for $\gamma<\frac{3}{2}$ according to eq.(58) is
$t_{f}=\frac{1}{3-2\gamma}G_{0}^{2}M_{0}^{3}\hbar^{-1}\alpha^{-1}.$ (64)
It is infinite for $\gamma\geq\frac{3}{2}$ but the mass is decreasing asymptotically to zero.
6 Summary
We have shown that the Bianchi identity of classical Einstein equations admits a violation of the fluid energy-momentum conservation if the Newton “constant” G and the cosmological term depend on time. The covariant energy-momentum divergence of the fluid appearing in the Bianchi identity has the form of the heat flow. We define the mass of the fluid as a volume integral of its energy density. If the pressure of the system is zero then the Bianchi identity relates $\partial_{t}(GM)$ to the time derivative of the cosmological term. For a non-zero pressure an analogous formula involves the temperature and entropy. A radiation of the fluid (involving an energy loss) can be included in Einstein equations if the the Newton “constant” or the cosmological term depend on time. The radiation is described by the Stefan-Boltzmann law. The Stefan-Boltzmann formula depends on the temperature and the area of the emitting surface. When the collapse of the star is described by pasting of Schwarzschild solution with the FLWR solution for a fluid with a certain density, then the resulting black hole should have the mass $M$ related to this density. We use the mass $M$ for a determination of the temperature and the area. The black hole mechanics derived in classical Einstein gravity gives a relation between $G$, $M$, the entropy $S$ and temperature $T$. In some effective theories of gravity when $G$ is time-dependent this relation may be more involved. From the Bianchi identity we obtain an equation relating the time derivatives of $G$, $\Lambda$ and $M$ to the temperature $T$ and entropy $S$. We discussed various formulas for $T$ and $S$ some of them directly generalizing the Bekenstein-Hawking expression and some (for $\gamma\neq 1$ and $\gamma\neq\frac{2}{3}$) quite hypothetical. These assumptions give a relation $G\simeq M^{-\gamma}$. When the expression $G(M)$ is substituted in the Stefan-Boltzmann formula for the mass decrease then it gives an evaporation law without free parameters. Depending on the value of $\gamma$ the black hole temperature and luminosity may either tend to infinity at finite time (black hole explosion), be constant during the whole black hole life time or even tend to zero in finite or infinite time. The latter option is interesting for the proposals of the dark matter as consisting of primordial black holes. The study of the possible black holes evaporation laws should give useful hints for their search in astronomical observations.
@@ 6 Summary
Acknowledgement A part of the results of this paper is contained in JH’s bachelor thesis. JH thanks Dr. Aleksander Kozak for supervising the thesis. ZH thanks Hytham Alwrekat for discussions on some topics of this paper.